Assume that ExxonMobil uses a standard cost system for each of its refineries. For the Houston refinery, the monthly fixed overhead budget is $8,000,000 for a planned output of 5,000,000 barrels. For September, the actual fixed cost was $8,750,000 for 5,100,000 barrels.

Required

a. Determine the fixed overhead budget variance.

b. If fixed overhead is applied on a per-barrel basis, determine the volume variance.

Provide formulas and an explanation.
Answer #1

Fixed overhead budget variance

= Actual fixed overhead - Budgeted

= 8,750,000 - 8,000,000

= 750,000 Unfavorable

Budgeted overhead rate per barrel = 8,000,000/5,000,000 = 1.60

Budgeted overhead for actual output = 1.60*5,100,000

= 8,160,000

Volume variance = 8,160,000-8,000,000

= 160,000 Favourable